# Dirichlet energy and focusing NLS condensates of minimal intensity

### v2

@@ 2.1 Boutroux Quadratic differentials of quasi-momentum type and finite-gap solutions
To close the circle of identifications we need to tie the discussion of finite-gap solutions with the main problem which is the object of Theorems 1.1, 1.2, 1.3. The square of any quasi-momentum ${\mathrm{d}}\mathbf{p}$ described in the section above, is, by construction, a BM (1.6). Viceversa the square-root of a BM is the quasi-momentum of a finite gap solution. The only subtlety is here in the identification of the bands.
@@
For any connectivity pattern $M$ the minimizer $\mathfrak{F}$ of the Dirichlet energy (1.4) over $\mathbb{K}_{E,M}$ (to be established in Theorems 1.1, 1.2) is the spectrum of a finite–gap solution of the fNLS equation. The average intensity (1.12) of this solution coincides with twice the Dirichlet energy of this minimizer.
@@
3 Generalized quasimomenta associated to compact sets
@@
is, in general, valid quasi-everywhere on the support of ${\mathrm{d}}\rho_{K}$.
@@
We can now use the first Green identity and the well known restoration formula for a measure from its potential [32] to derive [^4] For simplicity we assume sufficient smmothness (piece-wise smoothness) of $\partial\mathcal{K}$.
@@
where $dA$ is the standard Lebesque measure in ${\mathbb{R}}^{2}$.
@@ 3.1.1 The non simply connected case
Denote the stagnation points of Proposition 3.4 by $z_{1},\dots,z_{s}$ with multiplicies $m_{1},\dots,m_{s}$, $\sum_{j=1}^{s}m_{j}=k$.
From each $z_{j}$ we take all the arcs of steepest descent trajectories of $\Phi=\mathrm{Im}\,\mathcal{P}$ up to either another stagnation point or $\mathcal{K}\cup{\mathbb{R}}$. Denote by $\Sigma:=\bigcup_{d}\sigma_{d}$ the union of these arcs.
@@
Proof. This is the same as showing that $\mathcal{K}\cup\Sigma$ is a tree (in the sense of graph theory), with set $\mathcal{V}$ of $k+s+1$ vertices consisting of one node associated to each of the connected components $\mathcal{K}_{0},\dots\mathcal{K}_{k}$, and one for each stagnation point $z_{1},\dots,z_{s}$. The set of edges $\mathcal{E}$ consists of the $\sum_{j=1}^{s}(m_{j}+1)=s+k$ arcs in $\Sigma$.
@@ 4 Jenkins’ interception property and the Dirichlet energy
We say that a poly-continuum $\mathcal{K}$ has Jenkins’ interception property with respect to $\mathfrak{F}$ if $\mathcal{L}^{\perp}_{d}(z)\cap\mathcal{K}\neq\emptyset$ for all $z$ in the relative interior of every arc of $\mathfrak{F}$.
@@
Theorem 4.2 implies that a small deformation of $\mathfrak{F}$ in the dominant direction increases ${\mathcal{I}}(\mathfrak{F})$. Here “small” means a small and smooth displacement of a point $z$ in the interior of $\mathfrak{F}$ as well as a small variation in the normal direction ${\bf n}(z)$ to $\mathfrak{F}$. In fact, these deformations do not have to be small as long as the connectivity (the topology) of $\mathfrak{F}$ is preserved. The above arguments show that any $\mathfrak{F}$ with S-property is a local minimum of the Dirichlet energy functional ${\mathcal{I}}(\mathfrak{F})$. Moreover, it must be the global minimum of ${\mathcal{I}}(\mathfrak{F})$ in the subclass of compacts in ${\mathbb{K}}_{E}$ with the same connectivity as $\mathfrak{F}$, i.e., in the class of connectivity perserving deformations.
@@
From the asymptotics of the quasimomenta (3.9) it follows that both $\mathbf{p}(z),\mathcal{P}(z)$ are invertible for sufficiently large $|z|$.
The proof of Theorem 4.2 is preceeded by several lemmas. We start by denoting by $R_{L}$ the rectangle in the $\zeta=u+iv$–plane of area $2L^{2}$:
@@
Suppose that $\mathfrak{F}$ has the SS–property (2.18) at all points of $\mathfrak{F}^{\circ}$. Then we say that a compact set $\mathcal{K}$ possesses Jenkins’ interception property (relative to $\mathfrak{F}$) if for every $z\in\mathfrak{F}$ we have $\mathcal{L}^{\perp}(z)\cap\mathcal{K}\neq\emptyset$ for at least one orthogonal trajectories emanating from $z$, possibly at $z$ itself. We denote the class of these sets by $\mathbb{K}_{\mathfrak{F}}$ .
@@
We consider $Q{\mathrm{d}}z^{2}$, a quasimomentum type quadratic differential with simple poles at the points $e_{j}\in E$, $j=1,\dots,N$. Let $\mathfrak{F}\in{\mathbb{K}}_{E}$ be the Zakharov–Shabat spectrum of $Q{\mathrm{d}}z^{2}$ and let $M=M(\mathfrak{F})$ be the connectivity matrix of $\mathfrak{F}$. We want to prove that $\mathfrak{F}$ is a minimizer of $\mathcal{I}(\mathcal{K})$ among the subclass ${\mathbb{K}}_{E,M}\subset{\mathbb{K}}_{E}$ that have the same or greater connectivity as $\mathfrak{F}$. According to Theorem 4.10, it is sufficient to prove that any $\mathcal{K}\in{\mathbb{K}}_{E,M}$ has Jenkins’ interception property (Definition 4.8) with respect to $\mathfrak{F}$.
@@
We complete this section with last corollary of Theorem 4.10. Consider the RS $\mathfrak{R}_{N}$ defined by the branchpoints $E\cup\bar{E}$. Then the Zakharov–Shabat spectrum $\mathfrak{F}\in{\mathbb{K}}_{E}$ of the quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$ minimizes $\mathcal{I}(\mathcal{K})$ among all possible (Schwarz symmetrical) branchcuts of $\mathfrak{R}_{N}$. Before formulating this result we want to describe such collections of branchcuts.
@@
The unique minimizer $\mathfrak{F}$ of the Dirichlet energy in the class $\mathbb{L}_{E}\subset{\mathbb{K}}_{E}$ is the Zakharov–Shabat spectrum of the real normalized quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$, i.e., Zakharov–Shabat spectrum is the minimizing poly-continuum among all possible branchcuts of $\mathfrak{R}_{N}$:
@@
Our arguments now are similar to the proof of Theorem 1.3 presented above. Let $\mathcal{K}\in\mathbb{L}_{E}$ be a set of branch-cuts for $\cal R_{N}$ branched at $E\cup\overline{E}$. Let $z_{0}\in\mathfrak{F}$ be in relative interior of $\mathfrak{F}$ and let the regions $G_{1},G_{2}$ be defined as in the proof of Theorem 1.3 above. We have seen that $E\cap G_{1}$ has an odd number of points, from which we deduce that not all connected components of $\mathcal{K}$ that intersect $G_{1}$ can be entirely contained in $G_{1}$. Indeed, each such component can only contain an even number of points of $E$. Thus $\mathcal{K}$ must intersect the boundary of $G_{1}$ and hence at least one of $\mathcal{L}^{\perp}(z_{0})$. Now the statement follows from Theorem 4.10. $\blacksquare$
@@ 5 Hausdorff continuity of the Dirichlet energy
The Dirichlet energy $D_{\mathcal{K},\Phi}$ of $G$ is defined by any of the following formulae
$\pi D_{\mathcal{K},\Phi}=\pi\iint_{H^{+}}|{\rm grad}\,G|^{2}{\mathrm{d}}^{2}z=2\ \iint_{\partial\mathcal{K}\times\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w){\mathrm{d}}\rho_{\mathcal{K}}(z)=2\int_{\mathcal{K}}\mathrm{Im}\,\Phi(z){\mathrm{d}}\rho_{\mathcal{K}}(z).$ (5.6)
These equalities imply that each of the expressions is strictly positive, which can be written also as the positivity of the expression
@@
The energy $\mathcal{I}_{\Phi}(\mathcal{K})$ for the Problem 5.1 is given by
@@
In this formula the Green function is extended to $\mathbb{C}$$\times{\mathbb{C}}$ by Schwartz-symmetry.
@@ 6 Existence of the minimizer in ${\mathbb{K}}_{E,M}$
We now revert to the original problem with the external field given by $\Phi=z$. The goal of this section is to prove that, for any set of anchors $E$ and connectivity matrix $M$, the Dirichlet energy $\mathcal{I}(\mathcal{K})$ attains its minimum in the class ${\mathbb{K}}_{E,M}$. In view of Theorem 5.6 all that we need is to show that there exists a fixed rectangle $R_{E}\subset{{\mathbb{H}}}\cup{\mathbb{R}}$ and a minimizing sequence $\{\mathcal{K}_{n}\}\subset{\mathbb{K}}_{E,M}$, such that $\mathcal{K}_{n}\subset R_{E}$ for all $n\in{\mathbb{N}}$.
@@ 7 Schiffer variations and S-curves
$0=\mathrm{Re}\,\left(\iint\left(\frac{h(z)-\overline{h}(w)}{z-\overline{w}}-\frac{h(z)-h(w)}{z-w}-\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+2\int\Phi^{\prime}(z)h(z){\mathrm{d}}\mu(z)\right)$ (7.1)
@@ Appendix A Green functions
$G(z;\infty)\leq\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{C}}$ (A.4)
@@
$G(z;w)=\ln|\widetilde{\zeta}|\leq\sqrt{\frac{{\rm dist}(\widetilde{z},\widetilde{\mathcal{K}})}{\widetilde{C}}}$ (A.7)
@@
Let $\mathcal{K}\subset{{\mathbb{H}}}$ be a Dirichlet regular compact set, $c$ the minimal capacity of its components and $\mathcal{C}$ anoter disjoint compact. Then there is a constant $S>0$ such that
@@ Appendix B Maximal connectivity case: relation with Kuzmina’s Jenkins-Strebel differentials
Proof. Kuzmina proved in [22] (see corrections in [23], in particular Theorem 3 ibidem) a theorem about the existence and uniqueness of a meromorphic quadratic differential with prescribed number of annular domains and disk domains in correspondence with second order poles with negative biresidues. If we specialize the theorem and require no annular domains and no double poles then Kuzmina’s theorem guarantees that the critical graph is connected and the complement in $\mathbb{P}^{1}$ is a union of domains of type half-plane; the half planes abut poles of higher orders with prescribed singular expansion.
@@
so that we identify $T_{1}$ with the residue of the quadratic differential. For our application we need to set $T_{0}=1$, $T_{1}=0$, and ${\bf A}=E\cup\overline{E}$; in this case the uniqueness stated in the theorem implies that the critical graph $\Gamma$ is invariant under conjugation, $\Gamma=\overline{\Gamma}$. In particular the real axis must belong to the critical graph because there must be always one critical arc extending to the pole along the critical diractions (which in our case are $\arg(z)=0,\pi$). This proves the existence of the Zakharov–Shabat spectrum with simply connected complement in $\mathbb{H}$}}. $\blacksquare$

### v3

@@ 2.1 Boutroux Quadratic differentials of quasi-momentum type and finite-gap solutions
To close the circle of identifications we need to tie the discussion of finite-gap solutions with the main problem which is the object of Theorems 1.2, 1.3, 1.4. The square of any quasi-momentum ${\mathrm{d}}\mathbf{p}$ described in the section above, is, by construction, a BM (1.14). Vice versa the square-root of a BM is the quasi-momentum of a finite gap solution. The only subtlety is here in the identification of the bands.
@@
For any admissible connectivity pattern $M$ (Def. 1.1) the minimizer $\mathfrak{F}$ of the Dirichlet energy (1.12) over $\mathbb{K}_{E,M}$ (see Theorems 1.2, 1.3) is the spectrum of a finite–gap solution of the fNLS equation. The average intensity (1.20) of this solution coincides with twice the Dirichlet energy of this minimizer.
@@
3 Generalized quasi-momenta associated to compact sets
@@
is, in general, valid quasi-everywhere (i.e. up to sets of zero capacity) on the support of ${\mathrm{d}}\rho_{K}$.
@@
We can now use the first Green identity and the well known restoration formula for a measure from its potential [32] to derive [^4] For simplicity we assume sufficient smoothness (piece-wise smoothness) of $\partial\mathcal{K}$.
@@
where $dA$ is the standard Lebesgue measure in ${\mathbb{R}}^{2}$.
@@ 3.1.1 The non simply connected case
Denote the stagnation points of Proposition 3.4 by $z_{1},\dots,z_{s}$ with multiplicities $m_{1},\dots,m_{s}$, $\sum_{j=1}^{s}m_{j}=k$.
From each $z_{j}$ we take all the arcs of steepest descent trajectories of $\Phi=\mathrm{Im}\,\mathcal{P}$ up to either another stagnation point or $\mathcal{K}\cup{\mathbb{R}}$. Denote by
$\Sigma:=\bigcup_{d}\sigma_{d}$ (3.10)
the union of these arcs.
@@
Proof. Recall the description and definition of $\Sigma$ above (3.10). The statement is the same as showing that $\mathcal{K}\cup\Sigma$ is a tree (in the sense of graph theory), with set $\mathcal{V}$ of $k+s+1$ vertices consisting of one node associated to each of the connected components $\mathcal{K}_{0},\dots\mathcal{K}_{k}$, and one for each stagnation point $z_{1},\dots,z_{s}$. The set of edges $\mathcal{E}$ consists of the $\sum_{j=1}^{s}(m_{j}+1)=s+k$ arcs in $\Sigma$.
@@ 4 Jenkins’ interception property and the Dirichlet energy
We say that a poly-continuum $\mathcal{K}$ has Jenkins’ interception property with respect to $\mathfrak{F}$ if $\mathcal{L}^{\perp}_{d}(z)\cap\mathcal{K}\neq\emptyset$ for all $z$ in the relative interior, $\mathfrak{F}^{\circ}$, of every arc of $\mathfrak{F}$.
@@
Theorem 4.2 implies that a small deformation of $\mathfrak{F}$ in the dominant direction increases ${\mathcal{I}}(\mathfrak{F})$. Here “small” means a small and smooth displacement of a point $z$ in the interior of $\mathfrak{F}$ as well as a small variation in the normal direction ${\bf n}(z)$ to $\mathfrak{F}$. In fact, these deformations do not have to be small as long as the connectivity (the topology) of $\mathfrak{F}$ is preserved. The above arguments show that any $\mathfrak{F}$ with S-property is a local minimum of the Dirichlet energy functional ${\mathcal{I}}(\mathfrak{F})$. Moreover, it must be the global minimum of ${\mathcal{I}}(\mathfrak{F})$ in the subclass of compacts in ${\mathbb{K}}_{E}$ with the same connectivity as $\mathfrak{F}$, i.e., in the class of connectivity preserving deformations.
@@
From the asymptotics of the quasi-momenta (3.9) it follows that both $\mathbf{p}(z),\mathcal{P}(z)$ are invertible for sufficiently large $|z|$.
The proof of Theorem 4.2 is preceded by several lemmas. We start by denoting by $R_{L}$ the rectangle in the $\zeta=u+iv$–plane of area $2L^{2}$:
@@
Suppose that $\mathfrak{F}$ has the SS–property (2.18) at all points of $\mathfrak{F}^{\circ}$. We denote by $\mathbb{K}_{\mathfrak{F}}$ the class of the poly-continua that have Jenkins’ interception property relative to $\mathfrak{F}$ (Def. 4.1).
@@
We consider $Q{\mathrm{d}}z^{2}$, a quasimomentum type quadratic differential with simple poles at the points $e_{j}\in E$, $j=1,\dots,N$. Let $\mathfrak{F}=\mathfrak{F}_{Q}\in{\mathbb{K}}_{E}$ be the Zakharov–Shabat spectrum of $Q{\mathrm{d}}z^{2}$ and let $M(\mathfrak{F})$ be the connectivity matrix of $\mathfrak{F}$. We need to prove that $\mathfrak{F}$ is a minimizer of $\mathcal{I}(\mathcal{K})$ among the subclass ${\mathbb{K}}_{E,M(\mathfrak{F})}\subset{\mathbb{K}}_{E}$ that have the same or greater connectivity as $\mathfrak{F}$. According to Theorem 4.10, it is sufficient to prove that any $\mathcal{K}\in{\mathbb{K}}_{E,M(\mathfrak{F})}$ has Jenkins’ interception property (Definition 4.8) with respect to $\mathfrak{F}$, namely, that ${\mathbb{K}}_{E,M(\mathfrak{F})}={\mathbb{K}}_{\mathfrak{F}}$, the latter being the class in Def. 4.8.
@@
We complete this section with last corollary of Theorem 4.10. Consider the two–sheeted Riemann surface $\mathfrak{R}_{N}$ branched at $E\cup\bar{E}$. Then the Zakharov–Shabat spectrum $\mathfrak{F}\in{\mathbb{K}}_{E}$ of the quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$ minimizes $\mathcal{I}(\mathcal{K})$ among all possible (Schwarz symmetrical) branchcuts of $\mathfrak{R}_{N}$. Before formulating this result we want to describe such collections of branchcuts.
@@
The unique minimizer $\mathfrak{F}$ of the Dirichlet energy function, see (1.12), in the class $\mathbb{L}_{E}\subset{\mathbb{K}}_{E}$ is the Zakharov–Shabat spectrum of the real normalized quasimomentum differential ${\mathrm{d}}\mathbf{p}_{N}$ on $\mathfrak{R}_{N}$, i.e., Zakharov–Shabat spectrum is the minimizing poly-continuum among all possible branchcuts of $\mathfrak{R}_{N}$:
@@
Our arguments now are similar to the proof of Theorem 1.4 presented above. Let $\mathcal{K}\in\mathbb{L}_{E}$ be a set of branch-cuts for $\cal R_{N}$ branched at $E\cup\overline{E}$. Let $z_{0}\in\mathfrak{F}$ be in relative interior of $\mathfrak{F}$ and let the regions $G_{1},G_{2}$ be defined as in the proof of Theorem 1.4 above. We have seen that $E\cap G_{1}$ has an odd number of points, from which we deduce that not all connected components of $\mathcal{K}$ that intersect $G_{1}$ can be entirely contained in $G_{1}$. Indeed, each such component can only contain an even number of points of $E$. Thus $\mathcal{K}$ must intersect the boundary of $G_{1}$ and hence at least one of $\mathcal{L}_{d}^{\perp}(z_{0})$. Now the statement follows from Theorem 4.10. $\blacksquare$
@@ 5 Hausdorff continuity of the Dirichlet energy
The Dirichlet energy $\mathcal{I}_{\Phi}(\mathcal{K})$ of $G$ is defined by any of the following formulae
$\pi\mathcal{I}_{\Phi}(\mathcal{K})=\pi\iint_{H^{+}}|{\rm grad}\,G|^{2}{\mathrm{d}}^{2}z=2\ \iint_{\partial\mathcal{K}\times\partial\mathcal{K}}\ln\left|\frac{z-\overline{w}}{z-w}\right|{\mathrm{d}}\rho_{\mathcal{K}}(w){\mathrm{d}}\rho_{\mathcal{K}}(z)=2\int_{\mathcal{K}}\mathrm{Im}\,\Phi(z){\mathrm{d}}\rho_{\mathcal{K}}(z).$ (5.6)
These equalities imply that each of the expressions here above is strictly positive, which can be written also as the positivity of the expression
@@
The Dirichlet energy, $\mathcal{I}_{\Phi}(\mathcal{K})$ (5.6), of the solution $G$ of the Problem 5.1 is given by
@@
In this formula the Green function is extended to $\mathbb{C}$$\times{\mathbb{C}}$ by Schwartz-symmetry and the Wirtinger derivative is $\partial_{z}=\frac{1}{2}\left(\partial_{x}-i\partial_{y}\right)$.
@@ 6 Existence of the minimizer in ${\mathbb{K}}_{E,M}$
We now revert to the original problem with the external field given by $\Phi=z$. The goal of this section is to prove that, for any set of anchors $E$ and connectivity matrix $M$, the Dirichlet energy $\mathcal{I}(\mathcal{K})$ attains its minimum in the class ${\mathbb{K}}_{E,M}$, which was defined in (1.11). In view of Theorem 5.6 all that we need is to show that there exists a fixed rectangle $R_{E}\subset{{\mathbb{H}}}\cup{\mathbb{R}}$ and a minimizing sequence $\{\mathcal{K}_{n}\}\subset{\mathbb{K}}_{E,M}$, such that $\mathcal{K}_{n}\subset R_{E}$ for all $n\in{\mathbb{N}}$.
@@ 7 Schiffer variations and S-curves
$0=\mathrm{Re}\,\left(\iint\left(\frac{h(z)-\overline{h}(w)}{z-\overline{w}}-\frac{h(z)-h(w)}{z-w}\right){\mathrm{d}}\mu(z){\mathrm{d}}\mu(w)+2\int\Phi^{\prime}(z)h(z){\mathrm{d}}\mu(z)\right)$ (7.1)
@@ Appendix A Green functions
$G(z;\infty)\leq\sqrt{\frac{{\rm dist}(z,\mathcal{K})}{C}}.$ (A.4)
@@
$G(z;w)=\ln|\widetilde{\zeta}|\leq\sqrt{\frac{{\rm dist}(\widetilde{z},\widetilde{\mathcal{K}})}{\widetilde{C}}}$. (A.7)
@@
Let $\mathcal{K}\subset{{\mathbb{H}}}$ be a Dirichlet regular compact set, $c$ the minimal capacity of its components and $\mathcal{C}$ another disjoint compact. Then there is a constant $S>0$ such that
@@ Appendix B Maximal connectivity case: relation with Kuzmina’s Jenkins-Strebel differentials
Proof. Kuzmina proved in [22] (see corrections in [23], in particular Theorem 3 ibidem) a theorem about the existence and uniqueness of a meromorphic quadratic differential with prescribed number of annular domains and disk domains in correspondence with second order poles with negative bi-residues. If we specialize the theorem and require no annular domains and no double poles then Kuzmina’s theorem guarantees that the critical graph is connected and the complement in $\mathbb{P}^{1}$ is a union of domains of type half-plane; the half planes abut poles of higher orders with prescribed singular expansion.
@@
so that we identify $T_{1}$ with the residue of the quadratic differential. For our application we need to set $T_{0}=1$, $T_{1}=0$, and ${\bf A}=E\cup\overline{E}$; in this case the uniqueness stated in the theorem implies that the critical graph $\Gamma$ is invariant under conjugation, $\Gamma=\overline{\Gamma}$. In particular the real axis must belong to the critical graph because there must be always one critical arc extending to the pole along the critical directions (which in our case are $\arg(z)=0,\pi$). This proves the existence of the Zakharov–Shabat spectrum with simply connected complement in $\mathbb{H}$}}. $\blacksquare$